# Best Practices for Multi-Fidelity Bayesian Optimization in Materials and Molecular Research

### v1

@@ Results
We also include a negative example where the conditions of one of these benchmarks are changed to an unfavorable scenario and the associated MFBO performance is overridden. We change the LF conditions of the molecule polarizability benchmark ($\rho$ = 0.5 and $R^{2}$ = 0.49) to illustrate how $\Delta$ is reduced with respect to the previous case. Figure 6c shows the result for the optimization, where $\Delta$ is reduced from a maximum value of 0.67 to 0.03, providing almost no advantage over SFBO if compared with the previous favorable scenario (similar to the results described for the synthetic functions, see Results). The previous benchmarks show how the application of MFBO in suitable scenarios provides an advantage over standard SFBO cases and effectively decreases the optimization cost. By locating experiments where the cost and informativeness of the LF source are adequate, we can successfully translate the advantages of MFBO into a realistic experimental task.
@@ Discussion
Multi-fidelity Bayesian optimization (MFBO) offers a promising approach to reducing costs in optimization by leveraging inexpensive, approximate sources of information. In the chemical sciences, MFBO can accelerate optimization while maintaining cost efficiency. However, its utility is not universal; in some cases, it may prove less effective or even counterproductive18. To address these concerns, we conducted a comprehensive study to determine under which conditions MFBO outperforms standard single-fidelity Bayesian optimization (SFBO).
We introduce two key metrics for comparison: a standardized regret, $r$, and a discount metric, $\Delta$ (see Metrics). Our analysis of synthetic benchmarks demonstrates that MFBO performance is highly dependent on the characteristics of the low-fidelity (LF) source. Specifically, not all LF conditions are conducive to the effective use of MFBO. To systematically identify favorable conditions, we performed a grid-based exploration of LF cost and informativeness against the discount metric, $\Delta$. The results indicate that both $\rho$ and $R^{2}$ significantly influence the discount, with the most favorable outcomes achieved when the LF source is both highly informative and low-cost. Based on these findings, we developed a flowchart to guide the application of MFBO (see Figure 5). These guidelines were validated using three real-world experimental benchmarks in molecular and materials discovery, demonstrating that MFBO consistently outperforms SFBO in reducing optimization costs, provided the LF source is sufficiently inexpensive and informative.
Our work provides a structured decision-making framework for determining the applicability of MFBO. While our analysis focused on a specific surrogate model and two acquisition functions, the approach is extensible to other surrogate models and acquisition function families. Future research will expand this framework to include models such as Bayesian neural networks (BNNs)26, and explore additional experimental applications, evaluating the impact of varied feature spaces or additional fidelity levels. This study offers valuable insights for practitioners seeking to integrate MFBO into their experimental workflows, paving the way for more routine application of MFBO in chemical and materials optimization.
@@
The COFs benchmark data is available at https://github.com/SimonEnsemble/multi-fidelity-BO-of-COFs-for-Xe-Kr-seps., and it was extracted from 14. The polarizability dataset is available at https://zenodo.org/records/100471123. It was used in previous MFBO work13. The solvation energy dataset is extracted from 25. The code is available in this Github repository: https://github.com/atinary-technologies/chem-MFBO.git. It includes the instructions to run all the experiments described in this work.
@@ Appendix A Supplementary information
A.1 Synthetic functions expressions
@@
$f(\mathbf{x},l)=(x_{2}-(\frac{5.1}{4\pi^{2}}-0.1(1-\alpha))x^{2}_{1}+\frac{5}{\pi}x_{1}-6)^{2}+10(1-\frac{1}{8\pi})cos(x_{1})+10$ (2)
@@
$\begin{split}f(\mathbf{x},l)&=\frac{x_{1}}{2}(\sqrt{1+\frac{(x_{2}+x_{3}^{2})x_{4}}{x_{1}^{2}}}-1)\\ &+(x_{1}+(3-1.5(1-\alpha))x_{4})\exp(1+\sin(x_{3}))\end{split}$ (3)
@@
A.2 Simple regret algorithm
@@ 16             Append $r_{\rm{min}}$ to ${r}$
A.3 Informativeness measurement
The informativeness of the different problems is computed using the $R^{2}$ as explained in Metrics.

### v2

@@ Results
We also include a negative example where the conditions of one of these benchmarks are changed to an unfavorable scenario and the associated original MFBO performance is overridden. We artificially change the LF conditions of the molecule polarizability benchmark ($\rho$ = 0.5 and $R^{2}$ = 0.49) to illustrate how $\Delta$ is reduced with respect to the previous case. Figure 6c shows the result for the optimization, where $\Delta$ is reduced from a maximum value of 0.42 to 0.13 in the EI case (for MES, the performance decreases from 0.21 to -0.07), providing almost no advantage over SFBO if compared with the previous favorable scenario (similar to the results described for the synthetic functions, see Results). The previous benchmarks show how the application of MFBO in suitable scenarios provides an advantage over standard SFBO cases and effectively decreases the optimization cost. By locating experiments where the cost and informativeness of the LF source are adequate, we can successfully translate the advantages of MFBO into a realistic experimental task.
@@ Discussion
Multi-fidelity Bayesian optimization (MFBO) offers a promising approach to reducing costs in optimization by leveraging inexpensive, approximate sources of information. In the chemical sciences, MFBO can accelerate optimization while maintaining cost efficiency. However, its utility is not universal; in some cases, it may prove less effective or even counterproductive23. To address these concerns, we conducted a comprehensive study to determine under which conditions MFBO outperforms standard single-fidelity Bayesian optimization (SFBO).
Building on previous work in the field, we introduce two key metrics for comparison: a standardized regret, $r$, and a discount metric, $\Delta$ (see Metrics). Our analysis of synthetic benchmarks demonstrates that MFBO performance is highly dependent on the characteristics of the low-fidelity (LF) source. Specifically, not all LF conditions are conducive to the effective use of MFBO. To systematically identify favorable conditions, we performed a grid-based exploration of LF cost and informativeness against the discount metric, $\Delta$. The results indicate that both $\rho$ and $R^{2}$ significantly influence the discount, with the most favorable outcomes achieved when the LF source is both highly informative and low-cost. Based on these findings, we developed a flowchart to guide the application of MFBO (see Figure 5). These guidelines were validated using three real-world experimental benchmarks in molecular and materials discovery, demonstrating that MFBO consistently outperforms SFBO in reducing optimization costs, provided the LF source is sufficiently inexpensive and informative.
Our work provides a structured decision-making framework for determining the applicability of MFBO. While our analysis focused on a specific surrogate model and two acquisition functions, the approach is extensible to other surrogate models and acquisition function families. Future research will expand this framework to include models such as Bayesian neural networks (BNNs)33, and explore additional experimental applications, evaluating the impact of varied feature spaces or additional fidelity levels. This study offers valuable insights for practitioners seeking to integrate MFBO into their experimental workflows, paving the way for more routine application of MFBO in chemical and materials optimization.
@@
The COFs benchmark data is available at https://github.com/SimonEnsemble/multi-fidelity-BO-of-COFs-for-Xe-Kr-seps., and it was extracted from 16. The polarizability dataset is available at https://zenodo.org/records/100471130. It was used in previous MFBO work15. The solvation energy dataset is extracted from 29. The code is available in this Github repository: https://github.com/atinary-technologies/chem-MFBO.git. It includes the instructions to run all the experiments described in this work.
@@ Acknowledgements
Author contributions
V.S.G: conceptualization, methodology, software, investigation, visualization, writing - original draft. R.B.: conceptualization, methodology, software, writing - original draft. D.P.G.: conceptualization, methodology, software, writing - original draft. J.S.L: supervision, funding acquisition, resources, writing-review & editing. J.M.H-L: conceptualization, writing-review & editing. P.S.: supervision, funding acquisition, project administration, resources, writing-review & editing. L.R.: conceptualization, project administration, funding acquisition, supervision, resources, writing-review & editing.
@@ Appendix A Supplementary information
A.1 Preliminary investigation on the multi-fidelity kernel definition
We compare two different GP models that can be used in multi-fidelity tasks. Each model uses a specific kernel to model the covariance between fidelities. These kernels are part of the SingleTaskMultiFidelityGP and the MultiTaskGP models in BoTorch26 respectively, and they are used to model the fidelities of the samples.
The Downsampling kernel models the fidelity interaction in the SingleTaskMultiFidelityGP, and it has been already described in the Metrics section. The Index kernel models the fidelity interaction in the MultiTaskGP, and it is defined as:
$k(l,l^{\prime})=\left(BB^{\top}+\operatorname{diag}(\mathbf{\sigma})\right)_{l,l^{\prime}}$
if there are two fidelities and rank = 1 (current case), $B=\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$ and
$k(l,l^{\prime})=\begin{bmatrix}\alpha^{2}+\sigma&\alpha\beta\\ \alpha\beta&\beta^{2}+\sigma\end{bmatrix}.$
This results in an similar structure between the two fidelity kernels, where there are only three possible fidelity interactions, and the kernel scales the result of the input kernel depending on the corresponding interaction value.
We tested the performance of each model to compare the effect of the kernel. In a regression task, we sample 8 high-fidelity points in the Branin-2D function as a test set and train each surrogate with increasing training points (simulating the progressive acquisition of samples in the active learning loop). Models were trained with 10, 20, 30 and 40 samples (with a 50-50 ratio of high and low fidelity points) and the model performance was calculated using the $R^{2}$ computed on the 8 test set points. Figure 7 shows how both models have a similar performance in the high-data scenarios, with the default multi-fidelity kernel offering slightly better performance in the low-data scenario (10 training samples). For the MFBO case, the MT kernel was tested in the same settings as the favourable scenario of figure 3. Figure 8 shows how the performance of the model is similar to the default model, obtaining a slightly lower $\Delta$ (0.29).
Figure 7: Regression performance of each surrogate with increasing training data for the Branin task. MF: default multi-fidelity model with Downsampling kernel, MT: MultiTask GP model with Index kernel.
Figure 8: Results for BO loop using the MultiTask GP model with Index kernel and Expected Improvement on Branin-2D function under favourable conditions. The measured discount is $\Delta$=0.29
A.2 Synthetic functions expressions
@@
$f(\mathbf{x},\alpha)=(x_{2}-(\frac{5.1}{4\pi^{2}}-0.1(1-\alpha))x^{2}_{1}+\frac{5}{\pi}x_{1}-6)^{2}+10(1-\frac{1}{8\pi})cos(x_{1})+10$ (2)
@@
$\begin{split}f(\mathbf{x},\alpha)&=\frac{x_{1}}{2}(\sqrt{1+\frac{(x_{2}+x_{3}^{2})x_{4}}{x_{1}^{2}}}-1)\\ &+(x_{1}+(3-1.5(1-\alpha))x_{4})\exp(1+\sin(x_{3}))\end{split}$ (3)
@@
A.3 Simple regret algorithm
@@ 16             Append $r_{\rm{min}}$ to ${r}$
A.4 Discount diagrams
Figure 9 shows 3D-heatmaps of $\Delta$ as a function of $\rho$, $\alpha$ and $\tau$ for each synthetic function in the benchmark.
A.5 Informativeness measurement
The informativeness of the different problems is computed using the $R^{2}$ as explained in MetricsMetrics.